Graph (y+2)2 = -8(x-1)
Label vertex , focus, directrix

Answer by venugopalramana(3286)

(Y+2)^2=-8(X-1)
COMPARING WITH STD.EQN.OF PARABOLA
(Y-K)^2=4A(X-H)
WE GET
H=1
K=-2
A=-2
VERTEX=(H,K)=(1,-2)
FOCUS = (H+A,K)=(1-2,-2)=(-1,-2)
DIRECTRIX=X-H+A=0=X-1-2=0....OR....X=3